OpenStudy (anonymous):

The slope of a line is 2, and the y-intercept is 0. What is the equation of the line written in slope-intercept form? y = x + 2 y = 2 y = 2x

OpenStudy (whpalmer4):

Slope-intercept form is \(y = mx + b\) where \(m\) is the slope, and \(b\) is the y-intercept value. Plug in your values and see what you get :-)

OpenStudy (anonymous):

y = x + 2?

OpenStudy (whpalmer4):

The slope is 2, right? m = 2. Try again...

OpenStudy (anonymous):

so its y=2

OpenStudy (whpalmer4):

no.... \[y = mx + b\] \[m=2\]\[b=0\] what do you get?